OpenStudy (anonymous):

Before a nuclear reaction, there is 0.01000 kg of uranium. After the nuclear reaction, there is 0.00950 kg of uranium. How much energy was released? (h = 6.63 × 10^-34 J * s) A. 9.00 * 10^16 J B. 4.50 * 10^13 J C. 1.50 * 10^5 J D. 5.00 * 10^-4 J

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

here we have to apply the Einstein's formula: \[E = {c^2}\Delta m = 9 \times {10^{16}} \times \left( {0.01 - 0.0095} \right) = ...Joules\]

OpenStudy (anonymous):

ok! so we get 4.5E13! so our solution is choice B?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks! onto the next:)